OpenStudy (anonymous):

What is the simplified form of 1 over x squared all over 3 over x–cubed ?

OpenStudy (phi):

if the problem is \[ \frac{ \frac{1}{x^2}}{\frac{3}{x^3}} \] multiply top and bottom by x^3/3 (this changes the bottom fraction to 1 or think of it as : division is the same as multiply by the inverse (flipped version) either way: \[ \frac{1}{x^2} \cdot \frac{x^3}{3}\] can you finish?

OpenStudy (phi):

x^3 means x*x*x so this problems is \[ \frac{ x\cdot x \cdot x}{3\cdot x\cdot x} \] notice that x/x is 1, so you can "cancel" some of the x's

OpenStudy (anonymous):

is it 1/3x